aboyles:

If x represents the smallest integer, then which equation could be used to solve the problem? 2x + x + 4 = 25 2x + 2x + 4 = 25 2x + x + 2 = 25

UnluckySolaR:

Algebra 1?

aboyles:

yes

UnluckySolaR:

Ok, So do you have any clue or a guess?

aboyles:

i mainly have a guess but i need to get this right

UnluckySolaR:

What would you say it is?

aboyles:

i would say the 3rd one

UnluckySolaR:

Nope, But your close.

UnluckySolaR:

Its actually the 1st one 2x + x + 4 =25

aboyles:

oh thank you

UnluckySolaR:

Anytime Have a good one! :)

aboyles:

u 2

aboyles:

would you say this is right

UnluckySolaR:

Why do you think I told you the answer?

aboyles:

my friend grabbed my computer sorry

UnluckySolaR:

No worries.

surjithayer:

25-4=21 21=3*7 out of first two first is proper choice. for 3rd 25-2=23 it is a multiple by 1 and 23 so 3rd is not correct choice.

UnluckySolaR:

Thats what I said

UnluckySolaR:

ty for the medal

aboyles:

your welcome

aboyles:

hi 404